Each of the 10 firms in a competitive market has a



cost function of C = 25 + q2. The market demand function is Q = 120 - p. Determine the equilibrium



price, quantity per firm, and market quantity. M

1
Expert's answer
2021-10-31T18:29:02-0400

Solution:

Derive MC:

TC = 25 + q2


MC = "\frac{\partial TC} {\partial Q} = 2q"

TR = P "\times" Q

P = 120 – Q

TR = (120 – Q) Q = 120Q – Q2


MR = "\frac{\partial TR} {\partial Q} = 120 - 2q"

Set: MR = MC

120 – 2q = 2q

120 = 2q + 2q

120 = 4q

Q = 30

Equilibrium price = MC = 2q = 2 "\times" 30 = 60

Equilibrium price = 60

 

Market quantity: Q = 120 – P

Q = 120 – 60 = 60

Market quantity = 60

 

Quantity per firm = "\frac{60}{10} =" 6 units.